how do you solve the √1.25 and its solution?thanks=)
Answer by stanbon(75887)

how do you solve the √1.25 and its solution?
---
1.25 = 125/100 = (25)*5/100
------------------
So, sqrt(1.25)
= sqrt[(25*5)/100]
= (5/10)*sqrt5
= 0.5*sqrt(5)
======================================
Cheers,
Stan H.